# Uniform pathwise stability of additive singular SDEs driven by fractional Brownian motion

### v1

@@ 3.2 A priori estimates
Since $\Pi=\{t_{k}\}_{k=0}^{N}$ is a partition of $t_{0},t]$ such that $t_{k+1}-t_{k}\leq 1$ for $k\in\{0,\cdots,N\}$, we have
$|X_{t}-Y_{t}|^{2}$ $\leq e^{-2\kappa_{1}(t-t_{0})}|x-y|^{2}$
$\quad+2\sum_{k=0}^{N-1}e^{-2\kappa_{1}(t-t_{k+1})}\int_{t_{k}}^{t_{k+1}}e^{-2\kappa_{1}(t_{k+1}-r)}\langle X_{r}-Y_{r},b(X_{r})-b(Y_{r})\rangle dr.$
Therefore,
$\|X_{t}-Y_{t}\|_{L_{2m}}^{2}$ $\leq e^{-2\kappa_{1}(t-t_{0})}|x-y|^{2}$
$\quad+2\sum_{k=0}^{N-1}e^{-2\kappa_{1}(t-t_{k+1})}\|\int_{t_{k}}^{t_{k+1}}e^{-\kappa_{1}(t_{k+1}-r)}\langle X_{r}-Y_{r},b(X_{r})-b(Y_{r})\rangle dr\|_{L_{m}}.$
By applying Proposition 3.6 with $\psi=X-\widetilde{B}^{t_{0}}$, $\phi=Y-\widetilde{B}^{t_{0}}$ and using (3.34) , we get
$\begin{split}\|X_{t}-Y_{t}\|_{L_{2m}}^{2}&\leq e^{-2\kappa_{1}(t-t_{0})}|x-y|^{2}\\ &\quad+C\sum_{k=0}^{N-1}e^{-2\kappa_{1}(t-t_{k+1})}\|b\|_{\mathcal{C}^{\gamma}}(1+C_{0})\sup_{r\in[t_{k},t_{k+1}]}\|X_{r}-Y_{r}\|_{L_{2m}}^{2}(t_{k+1}-t_{k})^{1+H(\gamma-1)}\\ &\quad+C\sum_{k=0}^{N-1}e^{-2\kappa_{1}(t-t_{k+1})}\|b\|_{\mathcal{C}^{\gamma}}[X-Y]_{\mathcal{C}^{1/2}_{[t_{k},t_{k+1}]}L_{2m}}^{2}(t_{k+1}-t_{k})^{3/2+H(\gamma-1)}.\end{split}$
Let $\Pi=\{t_{k}\}_{k=0}^{N}$ be a partition of $t_{0},t]$ such that $t_{k+1}-t_{k}\leq 1$ for $k\in\{0,\cdots,N\}$. Using Lemma 3.11 and (3.34), we get
$\|X_{t}-Y_{t}\|_{L_{2m}}^{2}$ $\leq e^{-2\kappa_{1}(t-t_{0})}|x-y|^{2}+C(1+C_{0})^{2}\|b\|_{\mathcal{C}^{\gamma}}\sum_{k=0}^{N-1}e^{-2\kappa_{1}(t-t_{k+1})}\sup_{r\in[t_{k},t_{k+1}]}\|X_{r}-Y_{r}\|_{L_{2m}}^{2}.$ (3.36)
Denote by $g$ the function $g(t)=e^{2\kappa_{1}(t-t_{0})}\|X_{t}-Y_{t}\|_{L_{2m}}^{2}$, then we have
$g(t)\leq g(t_{0})+C\|b\|_{\mathcal{C}^{\gamma}}\sum_{k=0}^{N-1}\sup_{r\in[t_{k},t_{k+1}]}g(r).$
It follows that
$\sup_{t\in[t_{N-1},t_{N}]}g(t)\leq g(0)+C\|b\|_{\mathcal{C}^{\gamma}}\sum_{k=0}^{N-1}\sup_{r\in[t_{k},t_{k+1}]}g(r).$
@@
$g(t)\leq g(0)\exp\{C\|b\|_{\mathcal{C}^{\gamma}}(t-t_{0})\},$
which concludes the proof. ∎
@@
Hence using classical tightness arguments (e.g. [5, Theorem 7.3]), one can deduce that the sequence $(X^{n},W)_{n\in\mathbb{N}}$ is tight in $\mathcal{C}([t_{0},T])\times\mathcal{C}(\mathbb{R})$, where $\mathcal{C}([t_{0},T])$ the space of real continuous functions equipped with the topology of uniform convergence over compact sets and $\mathcal{C}(\mathbb{R})$ is the space of real continuous functions with sub-linear growth also equipped with the metric
$d(f,g)=\sum 2^{-k}\left(1\wedge\sup_{r\in[-2^{k},2^{k}]}|f(r)-g(r)|\right).$
Since this space is Polish, by the Prokhorov theorem, there exists a subsequence $(n_{k})_{k\in\mathbb{N}}$ such that $(X^{n_{k}},W)$ converges weakly in the space $\mathcal{C}([t_{0},T])\times\mathcal{C}(\mathbb{R})$.
@@ Conclusion.
The goal of this section is to prove Theorem 3.2. The proof is based on mollifying the drift, applying the regularisation properties and passing to the limit. Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered probability space and $(W_{t})_{t\in\mathbb{R}}$ be a two-sided $\mathbb{F}$-Wiener process. Solutions in this section will always be understood with respect to this probability space and $W$. Assume that $F$ satisfies Assumption 2.3. Let $\gamma\in(1-1/(2H),1)$ and $b$ satisfy Assumption 2.4. Let $x\in\mathbb{R}^{d}$, $T\in\mathbb{R}$, $t_{0}<T$, and $\left(b^{n}\right)_{n\in\mathbb{N}}$ be a sequence of bounded continuous functions converging to $b$ in $\mathcal{C}^{\gamma}$. Assume without any loss of generality that $\sup_{n\in\mathbb{N}}\|b^{n}\|_{\mathcal{C}^{\gamma}}<\Xi$.
@@
Since $X^{n}_{t_{0}}=X^{k}_{t_{0}}=x$, it follows that $\|\sup_{t\in[t_{0},T]}|X^{n}_{t}-X^{k}_{t}|\|_{L_{m}}$ converges to 0 as $n,k\rightarrow\infty$. Therefore, $(X^{n})_{n\in\mathbb{N}}$ is a Cauchy sequence in $L_{m}(\Omega;\mathcal{C}([t_{0},T]))$ and thus converges to a continuous process $X$. Since for any $n$, $X^{n}$ is adapted to $\mathbb{F}$, it follows that $X$ is also adapted to $\mathbb{F}$.
@@
Consequently, we deduce (again by Kolmogorov’s continuity theorem) that $(K^{n})_{n\in\mathbb{N}}$ is a Cauchy sequence in $L_{m}(\Omega,\mathcal{C}([t_{0},T]))$ and therefore converges to some continuous process $K$. In particular, $K$ satisfies (2.10) of Definition 2.1.
@@
By passing to the limit as $k\rightarrow\infty$ in (3.37), we get that $X^{n}-X]_{\mathcal{C}^{1/2}_{[t_{0},T]L_{m}}}\leq C\|b^{n}-b\|_{\mathcal{C}^{\gamma}}$. Hence, passing to the limit as $n\rightarrow\infty$ and using that $X^{n},K^{n}$ converge to $X,K$ in $L_{m}(\Omega,\mathcal{C}([t_{0},T]))$, we get that for any $t\in[t_{0},T]$, almost surely,
@@
4 Existence and uniqueness of invariant measures
The goal of this section is to prove existence and uniqueness of an invariant measure. Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a probability space and $W$ be an $\mathbb{F}$-Wiener process. Denote the left-sided Wiener measure by $\mathbf{W}$.
@@ 4 Existence and uniqueness of invariant measures
We follow an observation from [19] which considers evolutions of the SDE coupled with the historical past of the driving noise. To this end, we start the driving fBM from a deterministic continuous path $\bar{B}^{t_{0}}\equiv\eta$ and consider the evolution $t\mapsto\Psi(t_{0},t,x,\eta)$ given by the SDE
$\begin{split}d\Psi(t_{0},t,x,\eta)&=F(\Psi(t_{0},t,x,\eta))dt+b(\Psi(t_{0},t,x,\eta))dt+d\eta_{t}+d\widetilde{B}^{t_{0}}_{t},\,t\geq t_{0},\\ \Psi(t_{0},t_{0},x,\eta)&=x.\end{split}$ (4.1)
The driving signal consists of two parts: the past $\eta$ and the innovation $\widetilde{B}^{t_{0}}$. While it is not trivial that such SDE above has a unique solution over $t_{0},T]$, one can define a solution with respect to $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ and $W$ in a similar fashion to Definition 2.1 and prove well-posedness. In fact, for a smooth drift $b^{n}$, the above SDE has a unique solution $\Psi^{n}$ and the tightness and stability estimates presented in this paper still apply to $\Psi^{n}$ as they rely only on the regularisation properties of the innovation $\widetilde{B}^{t_{0}}$. So overall, one can obtain the same results by replacing $B$ by $\widetilde{B}^{t_{0}}$. These results are detailed in Theorem 4.2 and discussed further below. In the rest of the section, solutions will always be understood with respect to $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ and $W$.
A key point of [19] is that while the evolution (1.1) is not Markovian, the joint evolution of the solution and the increments of the noise is. To be more precise,
  1. 1.
    One defines the space $\mathcal{H}_{H}\subset\mathcal{C}_{0}((-\infty,0],\mathbb{R}^{d})$ to be the closure of $\mathcal{C}^{\infty}_{0}((-\infty,0])$ with respect to the norm
    $\|w\|_{\mathcal{H}_{H}}:=\sup_{s,t\in(-\infty,0]}\frac{|w(t)-w(s)|}{|t-s|^{(1-H)/2}\big(1+|t|+|s|\big)^{1/2}},$ (4.2)
    and shows that $\mathbf{W}$ is supported on this space (e.g. [19, Lemma 3.8]), and that the map $\mathbf{A}^{t_{0}}:\mathcal{H}_{H}\rightarrow\mathcal{C}([t_{0},\infty))$ defined by
    $\mathbf{A}^{t_{0}}:w\mapsto\Big(t\mapsto\alpha_{H}\int_{-\infty}^{t_{0}}\big((t-u)^{H-1/2}-(t_{0}-u)^{H-1/2}\big)dw_{u-t_{0}}\Big)$ (4.3)
    is continuous. Then one defines for any $w\in\mathcal{H}_{H}$ and $r\leq 0$, the process $\zeta(t_{0},t,w)(r):=\bar{\zeta}(t_{0},t,w)(r)-\bar{\zeta}(t_{0},t,w)(0)$ where
    $\bar{\zeta}(t_{0},t,w)(r)=\;\begin{cases}w\bigl(r+t-t_{0}\bigr),&r+t\leq t_{0},\\[6.0pt] W_{t+r}-W_{t_{0}},&t_{0}<r+t,\end{cases}$
@@
    This corresponds to concatenating an independent Wiener process to $w$ from 0 to $t-t_{0}$, re-parameterising the time parameter to $(-\infty,0]$, then re-centring the process in space to 0 at time 0. This ensures that the evolution of the noise stays in $\mathcal{H}_{H}\subset\mathcal{C}_{0}((-\infty,0],\mathbb{R}^{d})$, and that if $w\equiv(W(r+t_{0}))_{r\leq 0}$, then the resulting process $(\zeta(t_{0},t,w)(r))_{r\leq 0}$ has law $\mathbf{W}$.
  2. 2.
@@
    Finally, one can consider the Markovian evolution $\xi(t_{0},t,\cdot,\cdot)$ over $\mathbb{R}^{d}\times\mathcal{H}_{H}$ defined by
    $\xi(t_{0},t,x,w)=\left(\Psi(t_{0},t,x,\mathbf{A}^{t_{0}}(w)),\zeta(t_{0},t,w)\right).$ (4.4)
    Here it is understood that the process $(W_{t})_{t\geq t_{0}}$ used in the definition of $\zeta$ is the same one used for the evolution of $\Psi^{t_{0}}$ in the innovation of the fBm $\widetilde{B}^{t_{0}}$.
The semigroup associated with the evolution $\xi$ is then defined by
$(P_{t_{0},t}f)(x,w)=\mathbb{E}f(\xi(t_{0},t,x,w)),$ (4.5)
for any bounded measurable function $f:\mathbb{R}^{d}\times\mathcal{H}_{H}\rightarrow\mathbb{R}$, where the randomness is entirely given by $(W_{t})_{t\geq t_{0}}$.
@@
To define a meaningful concept of invariant measures, [19] first defines a generalised initial condition as a probability measure $\mu_{0}\in\mathcal{P}(\mathbb{R}^{d}\times\mathcal{H}_{H})$ such that its second marginal is equal to the left-sided Wiener measure. Then, $\mu$ is called an invariant measure associated to (4.1) if it is a generalised initial condition and it is invariant under the transition semigroup of $(\xi_{t})_{t\geq 0}$, that is $P_{t_{0},t}\mu=\mu$ for any $t\geq t_{0}$. Notice that the second marginal of the law of $\xi$ started from $\mu_{0}$ is always $\mathbf{W}$.
@@
In [19], the invariant measure is built within the more general stochastic dynamical systems (SDS) framework, where one constructs the SDS from the deterministic flow corresponding to the SDE (seen as a function of time, the initial condition and the driving signal). When $b$ is smooth, the first component of $\xi$ corresponds to the SDS and the second component corresponds to the stationary noise process. However, this approach is not well adapted when $b$ is singular since the solution of (4.1) is not given as a continuous function of the noise. We adopted a simpler construction here that is suited to our context. For completeness, we note that the literature also offers another construction via random dynamical systems (RDS). In fact, the evolution $\xi$ can be seen as an RDS if $b$ is smooth but this is difficult to check when $b$ is singular.
Remark 4.1.
The reason we consider the evolution of the noise over the space $\mathcal{C}((-\infty,0])$ and not $\mathcal{C}((-\infty,t_{0}])$ is to be able to compare invariant measures associated to SDEs started from different initial times. Let $t_{1}>t_{0}$ and $f:\mathbb{R}^{d}\times\mathcal{H}_{H}\rightarrow\mathbb{R}$ be a bounded measurable function. Then, using the definition of $\zeta$ and the cocycle property of the SDE (4.1), one can show that $P_{t_{0},t}f=P_{t_{1},t}(P_{t_{0},t_{1}}f)$. Therefore, if $\mu_{0}$ is an invariant measure associated to $(\Psi(t_{0},t,x,A^{t_{0}}(w)))_{t\geq t_{0}}$, then $\mu_{t_{0}}=P_{t_{0},t_{1}}\mu_{t_{0}}$ is also an invariant measure associated to $(\Psi(t_{1},t,x,A^{t_{1}}(w)))_{t\geq t_{1}}$. Finally, if the invariant measure is unique for $t_{0}$, then it is unique and the same for any $t_{1}$.
The invariant measure we define this way is specifically related to the SDE (4.1), it is therefore important to know how it is related to the original SDE (1.1). Below, we also show in Theorem 4.3 how we can identify the solution of (4.1) as the solution of (1.1) by plugging $\eta$ to be the history of the fBm. Basically, we will be able to write that for any $t$, almost surely, $\Psi(t_{0},t,x,\bar{B}^{t_{0}})=X_{t}$.
Main results
There are three main results in this Section. Theorem 4.2 states existence and uniqueness of a solution to (4.1) and give some regularity estimates. Theorem 4.3 establishes the link between the two SDEs (1.1) and (4.1). Finally, Theorem 4.4 gives existence and uniqueness of an invariant measure.

### v2

@@ 3.2 A priori estimates
Let $\delta>0$, $N=\lfloor t-t_{0}\rfloor/\delta$ and $\Pi=\{t_{k}\}_{k=0}^{N}$ be a partition such that $t_{k+1}-t_{k}=\delta$ for $k=0,\cdots,N-1$. Define $t_{N+1}=t$, we have for any $\tilde{t}\in[t_{N},t_{N+1}]$
$|X_{\tilde{t}}-Y_{\tilde{t}}|^{2}$ $\leq e^{-2\kappa_{1}(\tilde{t}-t_{0})}|x-y|^{2}$
$\quad+2\sum_{k=0}^{N-1}e^{-2\kappa_{1}(\tilde{t}-t_{k+1})}\int_{t_{k}}^{t_{k+1}}e^{-2\kappa_{1}(t_{k+1}-r)}\langle X_{r}-Y_{r},b(X_{r})-b(Y_{r})\rangle dr$
$\quad+\int_{t_{N}}^{\tilde{t}}e^{-2\kappa_{1}(\tilde{t}-r)}\langle X_{r}-Y_{r},b(X_{r})-b(Y_{r})\rangle dr$
By applying Proposition 3.6 with $\psi=X-\widetilde{B}^{t_{0}}$, $\phi=Y-\widetilde{B}^{t_{0}}$ and using (3.34), bounding $\sup_{[t_{N},\tilde{t}]}$ by $\sup_{[t_{N},t_{N+1}]}$, and $\tilde{t}-t_{N}$ by $t_{N+1}-t_{N}$ we get
$\begin{split}\|X_{\tilde{t}}-Y_{\tilde{t}}\|_{L_{2m}}^{2}&\leq e^{-2\kappa_{1}(\tilde{t}-t_{0})}|x-y|^{2}\\ &\quad+C\sum_{k=0}^{N}e^{-2\kappa_{1}(\tilde{t}-t_{k+1})}\|b\|_{\mathcal{C}^{\gamma}}(1+C_{0})\sup_{r\in[t_{k},t_{k+1}]}\|X_{r}-Y_{r}\|_{L_{2m}}^{2}(t_{k+1}-t_{k})^{1+H(\gamma-1)}\\ &\quad+C\sum_{k=0}^{N}e^{-2\kappa_{1}(\tilde{t}-t_{k+1})}\|b\|_{\mathcal{C}^{\gamma}}[X-Y]_{\mathcal{C}^{1/2}_{[t_{k},t_{k+1}]}L_{2m}}^{2}(t_{k+1}-t_{k})^{3/2+H(\gamma-1)}.\end{split}$
Using Lemma 3.11 and (3.34), we get
$\|X_{\tilde{t}}-Y_{\tilde{t}}\|_{L_{2m}}^{2}$ $\leq e^{-2\kappa_{1}(\tilde{t}-t_{0})}|x-y|^{2}+C(1+C_{0})^{2}\|b\|_{\mathcal{C}^{\gamma}}\sum_{k=0}^{N}e^{-2\kappa_{1}(\tilde{t}-t_{k+1})}\sup_{r\in[t_{k},t_{k+1}]}\|X_{r}-Y_{r}\|_{L_{2m}}^{2}\delta^{1+H(\gamma-1)}.$ (3.36)
Denote by $g$ the function $g(u)=e^{2\kappa_{1}(u-t_{0})}\|X_{u}-Y_{u}\|_{L_{2m}}^{2}$, then we have
[…]
$=g(t_{0})+C\delta^{1+H(\gamma-1)}\|b\|_{\mathcal{C}^{\gamma}}\sup_{r\in[t_{N},t_{N+1}]}g(r)+C\|b\|_{\mathcal{C}^{\gamma}}\sum_{k=0}^{N-1}\sup_{r\in[t_{k},t_{k+1}]}g(r).$
Taking the supremum over $\tilde{t}$ in $t_{N},t_{N+1}]$, we get
$\sup_{r\in[t_{N},t_{N+1}]}g(r)$ $\leq g(t_{0})+C\delta^{1+H(\gamma-1)}\Xi\sup_{r\in[t_{N},t_{N+1}]}g(r)+C\|b\|_{\mathcal{C}^{\gamma}}\sum_{k=0}^{N-1}\sup_{r\in[t_{k},t_{k+1}]}g(r).$
Choosing $\delta=(1/(2C\Xi))^{\frac{1}{1+H(\gamma-1)}}$, we have
$\sup_{t\in[t_{N},t_{N+1}]}g(t)\leq 2g(t_{0})+2C\|b\|_{\mathcal{C}^{\gamma}}\sum_{k=0}^{N-1}\sup_{r\in[t_{k},t_{k+1}]}g(r)$
@@
$g(t)\leq 2g(t_{0})\exp\Big\{2C\|b\|_{\mathcal{C}^{\gamma}}(N+1)\Big\}\leq 2g(t_{0})\exp\Big\{2C\delta^{-1}\|b\|_{\mathcal{C}^{\gamma}}(t-t_{0})\Big\},$
which concludes the proof. Note that the constant in front $g(t_{0})$ can be chosen arbitrarily close to $1$, and this changes the constant in the exponential. ∎
@@
Hence using classical tightness arguments (e.g. [5, Theorem 7.3]), one can deduce that the sequence $(X^{n},W)_{n\in\mathbb{N}}$ is tight in $\mathcal{C}([t_{0},T],\mathbb{R}^{d})\times\mathcal{C}(\mathbb{R},\mathbb{R}^{d})$. Since this space is Polish (recall the metric (2.4)), by the Prokhorov theorem, there exists a subsequence $(n_{k})_{k\in\mathbb{N}}$ such that $(X^{n_{k}},W)$ converges weakly in the space $\mathcal{C}([t_{0},T],\mathbb{R}^{d})\times\mathcal{C}(\mathbb{R},\mathbb{R}^{d})$.
@@ Conclusion.
The goal of this section is to prove Theorem 3.2. The proof is based on mollifying the drift, applying the regularisation properties and passing to the limit. Let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered probability space and $(W_{t})_{t\in\mathbb{R}}$ be a two-sided $\mathbb{F}$-Wiener process. Solutions in this section will always be understood with respect to this probability space and $W$. Assume that $F$ satisfies Assumption 2.4. Let $\gamma\in(1-1/(2H),1)$ and $b$ satisfy Assumption 2.2. Let $x\in\mathbb{R}^{d}$, $T\in\mathbb{R}$, $t_{0}<T$, and $\left(b^{n}\right)_{n\in\mathbb{N}}$ be a sequence of bounded continuous functions converging to $b$ in $\mathcal{C}^{\gamma}$. Assume without any loss of generality that $\sup_{n\in\mathbb{N}}\|b^{n}\|_{\mathcal{C}^{\gamma}}\leq\Xi$.
@@
Since $X^{n}_{t_{0}}=X^{k}_{t_{0}}=x$, it follows that $\|\sup_{t\in[t_{0},T]}|X^{n}_{t}-X^{k}_{t}|\|_{L_{m}}$ converges to 0 as $n,k\rightarrow\infty$. Therefore, $(X^{n})_{n\in\mathbb{N}}$ is a Cauchy sequence in $L_{m}(\Omega;\mathcal{C}([t_{0},T],\mathbb{R}^{d}))$ and thus converges to a continuous process $X$. Since for any $n$, $X^{n}$ is adapted to $\mathbb{F}$, it follows that $X$ is also adapted to $\mathbb{F}$.
@@
Consequently, we deduce (again by Kolmogorov’s continuity theorem) that $(K^{n})_{n\in\mathbb{N}}$ is a Cauchy sequence in $L_{m}(\Omega,\mathcal{C}([t_{0},T],\mathbb{R}^{d}))$ and therefore converges to some continuous process $K$. In particular, $K$ satisfies (2.11) of Definition 2.1.
@@
By passing to the limit as $k\rightarrow\infty$ in (3.37), we get that $X^{n}-X]_{\mathcal{C}^{1/2}_{[t_{0},T]L_{m}}}\leq C\|b^{n}-b\|_{\mathcal{C}^{\gamma}}$. Hence, passing to the limit as $n\rightarrow\infty$ and using that $X^{n},K^{n}$ converge to $X,K$ in $L_{m}(\Omega,\mathcal{C}([t_{0},T],\mathbb{R}^{d}))$, we get that for any $t\in[t_{0},T]$, almost surely,
@@
4 Markovian structure and invariant measures
The goal of this section is to prove the existence and uniqueness of an invariant measure. We fix a probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ and an $\mathbb{F}$-Wiener process $W$. Let $\mathcal{C}_{0}((-\infty,0],\mathbb{R}^{d})$ (respectively $\mathcal{C}^{\infty}_{0}((-\infty,0],\mathbb{R}^{d})$) denote the space of $\mathbb{R}^{d}$-valued continuous (respectively smooth) functions over $(-\infty,0]$ that vanish at 0. Let $\mathbf{W}$ be the left-sided Wiener measure, i.e. the measure induced by $W|_{(-\infty,0]}$ on $\mathcal{C}_{0}((-\infty,0],\mathbb{R}^{d})$. Let $\mathcal{H}_{H}$ be the closure of $\mathcal{C}^{\infty}_{0}((-\infty,0],{\mathbb{R}^{d}})$ with respect to the norm
$\|w\|_{\mathcal{H}_{H}}:=\sup_{s,t\in(-\infty,0]}\frac{|w(t)-w(s)|}{|t-s|^{(1-H)/2}\big(1+|t|+|s|\big)^{1/2}}.$ (4.1)
Then, $\mathcal{H}_{H}$ is Polish and $\mathbf{W}$ is supported on $\mathcal{H}_{H}$ ([21, Lemma 3.10 and Lemma 3.8]).
@@ 4 Markovian structure and invariant measures
The first enhancement is to allow dynamics of (1.1) to start from a given memory. To this end, it is necessary to replace the driving signal up to time $t_{0}$, i.e. $\bar{B}^{t_{0}}$, by a generic $\mathcal{F}_{t_{0}}$-measurable continuous path $V$ and consider the corresponding SDE to (1.1) starting from $(X_{0},V)$ at time $t_{0}$. More precisely, we consider the evolution $t\mapsto X^{(X_{0},V)}(t_{0},t)$ given by the SDE
$\begin{split}dX^{(X_{0},V)}(t_{0},t)&=F(X^{(X_{0},V)}(t_{0},t))dt+b(X^{(X_{0},V)}(t_{0},t))dt+dV_{t}+d\widetilde{B}^{t_{0}}_{t},\quad t\geq t_{0},\\ X^{(X_{0},V)}(t_{0},t_{0})&=X_{0}.\end{split}$ (4.2)
The initial input at time $t_{0}$ consists of $X_{0}$ and $V$ which are $\mathcal{F}_{t_{0}}$-random variables in $\mathbb{R}^{d}$} and $\mathcal{C}_{0}([0,\infty),{\mathbb{R}^{d}})$ respectively. The innovation $\widetilde{B}^{t_{0}}$ plays the role of driving signal. Given $t_{0},X_{0},V$, one can define a solution with respect to $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ and $W$ in a similar fashion to Definition 2.1. Earlier results for (1.1) rely solely on the regularisation properties of the innovation $\widetilde{B}^{t_{0}}$, and therefore, can be carried forward to equation (4.2), see Theorem 4.1 below. In particular, when $x\in{\mathbb{R}^{d}}$ and $V=\bar{B}^{t_{0}}$, (4.2) becomes (1.1), and hence, by uniqueness, $(X^{(x,\bar{B}^{t_{0}})}(t_{0},t))_{t\geq 0}$ is the solution to (1.1) starting at $x$. It will also be useful to consider the evolution from a deterministic initial input $(x,v)$. In Theorem 4.3, we show that there exists a jointly measurable map
$(t_{0},t,x,v)\mapsto X(t_{0},t,x,v),$ (4.3)
such that for any $t\geq t_{0}$, almost surely in $\omega$, $X(t_{0},t,x,v)=X^{(x,v)}(t_{0},t)$. We also prove that for any $\mathcal{F}_{t_{0}}$-random variable $(X_{0},V)$ and $t\geq t_{0}$, we have $X(t_{0},t,X_{0}(\omega),V(\omega))(\omega)=X^{(X_{0},V)}(t_{0},t)(\omega)$ almost surely in $\omega$.
The map (4.3) will be important in defining a Markov evolution. In fact, a key point of [21] is that while the evolution (1.1) is not Markovian, the joint evolution of (4.3) and the increments of the noise is. The evolution of the noise component takes place in the state space $\mathcal{H}_{H}\subset\mathcal{C}_{0}((-\infty,0],\mathbb{R}^{d})$. Then one defines for any $w\in\mathcal{H}_{H}$ and $r\leq 0$, the process
$Z(t_{0},t,w)(r):=\bar{Z}(t_{0},t,w)(r)-\bar{Z}(t_{0},t,w)(0),$ (4.4)
where
$\bar{Z}(t_{0},t,w)(r)=\;\begin{cases}w\bigl(r+t-t_{0}\bigr),&\text{ when }r+t\leq t_{0},\\[6.0pt] W_{r+t}-W_{t_{0}},&\text{ when }t_{0}<r+t,\end{cases}$
@@
Heuristically, this corresponds to concatenating the Wiener path $(W_{r}-W_{t_{0}})_{r\geq t_{0}}$ to $(w({r-t_{0}}))_{r\leq t_{0}}$, and, to ensure that resulting process stays in $\mathcal{H}_{H}\subset\mathcal{C}_{0}((-\infty,0],\mathbb{R}^{d})$, re-parametrising the time parameter to $(-\infty,0]$, then re-centring the process in space to 0 at time 0. If one ignores the time re-parametrisation, then for each $t\geq t_{0}$, $Z(t_{0},t,w)$ corresponds to the entire trajectory up to time $t$ of the resulted centred concatenated path. Note that if $w\equiv(W_{r+t_{0}}-W_{t_{0}})_{r\leq 0}$, then the resulting process $(Z(t_{0},t,w)(r))_{r\leq 0}$ is nothing but $(W_{t+r}-W_{t})_{r\leq 0}$, which has law $\mathbf{W}$.
To couple $Z(t_{0},t,w)$ with (4.3), we use the continuous map (see [21, Lemma 3.8])
$\mathbf{A}^{t_{0}}:\mathcal{H}_{H}$ $\rightarrow\mathcal{C}([t_{0},\infty),{\mathbb{R}^{d}})$ (4.5)
$w$ $\mapsto\Big(t\mapsto\alpha_{H}\int_{-\infty}^{t_{0}}\big((t-u)^{H-1/2}-(t_{0}-u)^{H-1/2}\big)dw(u-t_{0})\Big),$
@@
and consider the jointly measurable map $\Theta(t_{0},t,\cdot,\cdot)$ over $\mathbb{R}^{d}\times\mathcal{H}_{H}$ defined by
$\Theta(t_{0},t,x,w)=\left(X(t_{0},t,x,\mathbf{A}^{t_{0}}(w)),Z(t_{0},t,w)\right)$ (4.6)
and for each $s\leq t$, the corresponding operator
$(P_{s,t}f)(x,w)=\mathbb{E}f(\Theta(s,t,x,w)),$ (4.7)
for any bounded measurable function $f:\mathbb{R}^{d}\times\mathcal{H}_{H}\rightarrow\mathbb{R}$.
@@
To connect the evolution $\Theta$ with (1.1), [21] defines a generalised initial condition as a probability measure $\mu_{0}\in\mathcal{P}(\mathbb{R}^{d}\times\mathcal{H}_{H})$ such that its second marginal is the left-sided Wiener measure $\mathbf{W}$. It turns out (Theorem 4.3) that the evolution $\Theta$ is a Markov process with evolution family $P$ (see Definition 4.7), which becomes time-homogeneous when started from a generalised initial condition. Furthermore, because the second marginal of the law of $\Theta$ started from a generalised initial condition is always $\mathbf{W}$, the set of generalised initial conditions is invariant under $(P_{s,t})$. Hence, defining $P_{t}=P_{0,t}$, then $(P_{t})_{t\geq 0}$ forms a transition semigroup on the set of generalised initial conditions. As in [21], we call $\mu$ an invariant measure (see Definition 4.4) associated to (4.2) if it is a generalised initial condition and it is invariant under the transition semigroup $(P_{t})$, that is $P^{*}_{t}\mu=\mu$ for any $t\geq 0$ where $P_{t}^{*}$ is the adjoint operator.
@@
In [21], the invariant measure is built within the more general stochastic dynamical systems (SDS) framework, where one constructs the SDS from the deterministic flow corresponding to the SDE (seen as a function of time, the initial condition and the driving signal). When $b$ is smooth (so that (4.2) can be treated as a random ordinary differential equation), the first component of $\Theta$ corresponds to the SDS and the second component corresponds to the stationary noise process. However, this approach is not well adapted when $b$ is singular since the solution of (4.2) is not given as a continuous function of the driving noise. We have adopted a simpler construction, relying solely on the theory of Markov processes, which is well suited to our context.
Main results.